OpenStudy (anonymous):

find the point on the graph of f(x)=sqrt(x-8) that is closest to the point (12,0)

OpenStudy (anonymous):

Let \((a,b)=(a,\sqrt{a-8})\) be the point on the curve. The distance between \((a,b)\) and \((12,0)\) will be \[\begin{align*}d&=\sqrt{(a-12)^2+(b-0)^2}\\ &=\sqrt{(a-12)^2+(\sqrt{a-8})^2}\\ &=\sqrt{a^2-24a+144+(a-8)}\\ &=\sqrt{a^2-23a+136} \end{align*}\] Use the first derivative test to figure out the \(a\) that minimizes \(d\). One thing to note is that, given \(f(x)=\sqrt{g(x)}\), \(f\) is optimized for the same \(x\) as \(g\) is, which means you can use the derivative test on \[d^*=a^2-23a+136\] to solve the problem.